OpenStudy (anonymous):

George has found a relationship between the number of people at an amusement park in the afternoon and the air temperature, in degrees Celsius. The table below shows the data collected by George: Temperature (in degrees Celsius) (x) 20 21 22 23 24 25 26 27 28 Number of people at the amusement park in the afternoon (y) 200 190 181 172 160 150 140 130 120 Part A: What would most likely be the number of people at the amusement park in the afternoon if the temperature is 35 degrees Celsius? (3 points) Part B: Predict a possible correlation coefficient for the data in the table and explain why

jigglypuff314 (jigglypuff314):

Hello :) I saw you ask for help in chat, I'm not very good at word problems but I'll try to help ^_^ so every time the temperature increases by 1, the number of people seems to decrease by about 10 do you see this? :)

OpenStudy (anonymous):

yes:)

jigglypuff314 (jigglypuff314):

so if you keep following that pattern... 28 29 30 31 120 110 100 90 etc what do you think you would get at 35?

OpenStudy (anonymous):

i think part a would be 50 but i dont get part b or c

jigglypuff314 (jigglypuff314):

great :) A is correct ^_^ I don't see a C mmm B is asking for slope, do you know slope? :3

OpenStudy (anonymous):

part c is Part C: George says that if a discount is given on the price of tickets at the amusement park, the number of people at the park would increase. Is this an example of correlation or causation? Justify your answer. and i kind of do i have geogebra that i think i can make one

jigglypuff314 (jigglypuff314):

ah I do not know the terms "correlation" and "causation" sorry :/ \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]when given points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

thank you but im still a lil confused lol

jigglypuff314 (jigglypuff314):

when I said "every time the temperature increases by 1, the number of people seems to decrease by about 10" that is basically slope slope = rise / run rise = decrease 10 run = temperature increase 1 so slope = -10/1 = ?